### v2

@@ 4.1 Different Prompting Strategies
Our results applying these strategies to Mixtral 8x7B Instruct v1 are included in Table 2. We also provide comprehensive results in Appendix B.2 including for the LLAMA-2 70B Chat and LLAMA-3 70B Instruct models (Table 7) and Mistral Large 2 (Table 8). We find that LM Prompting always achieves the best literal theory of mind performance. Adding the QA format seems to make literal theory of mind a lot worse by taking the input distribution further from the LLM task and closer to the kind of processing needed for actual decision making. The gap between functional theory of mind and literal theory of mind performance is predictably smaller in this case. CoT Prompting has an inconsistent effect on literal theory of mind and always appears to make the functional theory of mind performance better than vanilla QA Prompting. Reflexion and Plans+Insights also seem to improve on QA prompting in terms of functional theory of mind performance, although we find that they do not quite match the performance of generic CoT prompting.
@@ 4.2 Reasoning Over Long Contexts
Deeper Analysis. We have conducted a number of experiments that we do not have space to recap in detail within the main text. In Appendix B.3 we validate the generality of our findings regarding functional theory of mind performance across open source models including LLAMA-3 70B Instruct, Mixtral 8x7B Instruct v1, and Mistral Large 2 when paired with more dynamic tit-for-tat style partner policies. Additionally, in Appendix B.4 we look at the role of inductive bias in the action representation on performance of LLAMA-3 70B Instruct and Mistral Large 2 models. We find that more inductive bias is helpful with a small number of interactions, but could lead to poor long-term behavior that does not converge with respect to the partner’s policy.
@@ 5 Alternative Views
We should embrace game theory, not theory of mind. This is perhaps the most prominent alternative view to our paper that was already highlighted in our Rock, Paper, Scissors example. Many papers in the multi-agent RL literature consider finding a game theoretic equilibrium solution as their ultimate goal (Littman et al., 2001; Wang & Sandholm, 2002; Greenwald et al., 2003; Zinkevich et al., 2005; 2007). In some ways game theoretic solutions are opposite to solutions that are found with theory of mind because they focus on performance against worst case or optimal partners rather than the partners agents they are actually paired with. Likewise, they focus on eliminating the exploitability of solutions by other agents as opposed to maximizing expected return in the context of the actual policies of these agents. Proponents of the game theoretic approach will readily acknowledge that theory of mind is more descriptive of the way that humans tend to act (Colman, 2003; Larson, 2004; Owen, 2013). The argument is rather that game theory is prescriptive of the way that AI agents should ideally act. We believe that there is indeed some validity to this point and acknowledge that in a closed system in which all involved agents are AI, this is the optimal case. However, this does not hold as the community moves towards increasingly real-world use cases in which AI must interact with humans or AI agents that behave sub-optimally. These use cases are the norm for LLMs and are the setting in which researchers are interested in measuring LLM theory of mind capabilities. Here there is a gap between the expected reward achievable with the theory of mind solution that optimizes for the expected case and the game theoretic solution that optimizes for the worst case. Moreover, this gap should only get larger as the agent interacts with more humans. We definitely acknowledge that what society will end up preferring depends on the alignment of incentive structures. Humans will probably prefer theory of mind solutions when their relationship with the agent is mostly collaborative, but will not want to be exploited when their relationship with the agent is mostly adversarial. We strongly support instituting guardrails in these cases and believe that this makes it even more important to measure functional theory of mind performance to understand the extent of exploitation by AI models. We provide further detail about the connection between functional theory of mind and relevant equilibria concepts in Appendix A.
@@
Literal theory of mind is paramount for insights related applications. For some applications, we only desire that AI agents provide us with insights, which we will use to make decisions. For these applications, we do not wish to provide the AI with the agency to make these decisions for us. There are indeed very important use cases of this type for which literal theory of mind would be a sufficient capability without functional theory of mind. For example, providing theory of mind insights to humans that struggle with theory of mind (such as those with autism (Baron-Cohen, 1989; Perner et al., 1989; Tager-Flusberg, 1992; Fletcher-Watson et al., 2014)) or providing insights about the future behavior of foreign governments (Ye et al., 2024). Generally, these applications are most helpful when the AI has access to more information than the human pertaining to the situation. However, often times the amount of agency we decide to give to an AI depends on its expected reliability in making actual decisions. In this paper we argue that functional theory of mind serves as a necessary test for deciding if AI is ready to receive this agency and that literal theory of mind on its own is not reliable for assessing this performance. It is also important to note that in most applications where we do not provide AI with the agency to act, it would still be useful to humans to have the AI provide an action recommendation.
@@
The word ”broken” is too strong. It can be argued that because there are many ways to measure literal theory of mind, adding functional theory of mind is a just single additional aspect within a comprehensive theory of mind evaluation. However, we think ”broken” is an appropriate term here because functional theory of mind is what we really care about for applications where we give the AI agency to act, and we have found that deploying an LLM in this setting simply based on its literal theory of mind performance can be misleading to a potentially dangerous extent depending on the application. It has also been argued that process consistency will simply emerge with larger and more capable models (Bubeck et al., 2023). The position of our paper is agnostic to the method by which process consistency is achieved. Regardless, given that models like GPT-4 do not display process consistency (Bubeck et al., 2023), we believe it is of paramount importance to include measurements of functional theory of mind to directly measure the behavioral implications of the achieved level of process consistency in the context of assessing theory of mind performance.
@@
Functional theory of mind is easy with fine-tuning. RL style fine-tuning would easily solve the functional theory of mind problem displayed in our matrix game based experiments as highlighted with simple tabular models. However, our experiments still demonstrate a case where LLM literal theory of mind capabilities do not imply functional theory of mind capabilities and thus highlight the importance of directly measuring functional theory of mind. Most prominent LLM based services today also do not actually tune the model weights separately for each user and rely on in-context learning. The likely reason for this is partly about hardware related restrictions and partly about the difficulty of performing reliable continual learning. See Appendix A for a detailed discussion of connections to continual learning.
@@ 6 Conclusion
In this paper, we have argued that existing theory of mind benchmarks are broken because they measure literal theory of mind (Definition 2.1), and not functional theory of mind (Definition 2.2), which is what we typically actually care about in multi-agent social contexts. In Section 4 and Appendix B we provided comprehensive evidence that functional theory of mind performance can be quite poor even when literal theory of mind performance is good. Our analysis considers LLMs playing simple matrix games with very simple partners to underscore this point. This position paper serves as a call to action to the community to develop suitably complex LLM theory of mind benchmarks that directly measure functional theory of mind performance when acting with partners from different personas and social contexts.
Impact Statement
Our work aims to improve the theory of mind capabilities of AI agents by arguing that those capabilities depend both on making accurate predictions and acting upon those predictions in a rational way. If our viewpoint is widely adopted, AI agents will materially improve in their ability to respond to human intention and behavior. At a societal level, these improvements would be beneficial, such as through improved productivity from having more efficient AI agent user experiences. However, these improvements may also be harmful if the AI’s theory of mind is used to manipulate human behavior in nefarious, malicious, or undisclosed ways.
@@
Appendix A Connections to Equilibria Concepts, Continual Learning, and Potential Applications
@@
In-Context Learning vs. Continual Learning. In a multi-agent environment, the environment is nonstationary from the perspective of each agent if the policies of the other agents it interacts with are changing or learning over time (Littman, 1994). As a result, the setting of rapid adaptation to new agents in our paper can be considered a special case of continual RL (see Proposition 3 of Khetarpal et al. (2020)). Proposition 2 of Khetarpal et al. (2020) established that all continual RL problems can be modeled as partial observable problems with agents conditioned on the full interaction history. So, in principle, in-context learning over sufficiently long context representations with sufficiently expressive neural networks can be considered a general solution to continual learning problems. However, it remains to be seen if it is possible to see in-context behaviors resembling continual learning strategies like reservoir sampling replay buffers (Riemer et al., 2018a) or scalable memory efficient approximate buffers (Riemer et al., 2017a; 2019; Bashivan et al., 2019). It would be interesting to also consider policy changes from step to step and their degree of correspondence with older models similar to work leveraging knowledge distillation for continual learning (Li & Hoiem, 2017; Riemer et al., 2017b; Kirkpatrick et al., 2017). Intuitively, chain of thought reasoning may be beneficial because of its similarity with continual learning approaches that can select which layers to process at inference time (Rosenbaum et al., 2018; Cases et al., 2019; Chang et al., 2019; Rosenbaum et al., 2019a), see (Rosenbaum et al., 2019b) for a survey of approaches. More generally, this adaptive computation problem can be formalized within the coagent networks framework (Thomas, 2011; Kostas et al., 2020; Zini et al., 2020). That said, the compositional generalization needed for utilizing these models in practice makes achieving real-world success very challenging (Klinger et al., 2020), potentially making LLM style training a more viable strategy for learning this composition. That said, a word of caution comes from recent theoretical insights about the difficulty of efficiently evaluating models with large context lengths (Riemer et al., 2024a) acting in complex environments (Riemer et al., 2022). This could be a leading reason for the poor performance we are currently seeing with LLMs. Ultimately, there are many potentially interesting continual learning settings that should be considered for a comprehensive evaluation (Normandin et al., 2021), which we leave to future work.
@@
Alternatives to Continual Training. While our work highlights the limitations of using in-context learning for continual learning with current LLMs, full fine-tuning or continual training can be quite expensive in the context of LLMs – especially when it must be done for each end user. However, more computationally efficient alternatives like parameter efficient tuning (Hu et al., 2021; Dettmers et al., 2024; Thakkar et al., 2024a) or model merging (Ilharco et al., 2022; Yadav et al., 2024; Akiba et al., 2024; Thakkar et al., 2024b) may constitute a more economical middle ground. We leave exploration of the comparative efficacy of these strategies to future work. It is important to note that neural scaling laws only demonstrate generalization improvements with bigger model sizes when there is access to as much data as needed (Kaplan et al., 2020). In fact, capacity limits have been found to enhance generalization for RL (Malloy et al., 2020b; 2023) and multi-agent RL (Malloy et al., 2020a; 2021b; 2021a) in the limited data regime. Smaller models also may be necessary to maintain performance in realtime environments that are sufficiently stochastic (Riemer et al., 2024c; b). As such, it may make more sense to consider smaller LLMs if we plan on doing user specific finetuning or auxiliary objectives we would like to maintain like moral values (Padhi et al., 2024; Dognin et al., 2024).
@@
Appendix B Experiment Details
For our experiments we ran each LLM model using the Watsonx API playing a sequence of 100 step episodes with selected policies for the other agents over the course of 24 hours or a maximum of 100 episodes. Every model was evaluated for at least 30 episodes and confidence intervals are based on the actual sample size considered for each LLM model. The payoff tables used for each game are provided in Tables 4, 5, and 6. We also considered the following representations for each action as indicated in the main text:

### v3

@@ 4.1 Different Prompting Strategies
Our results applying these strategies to Mixtral 8x7B Instruct v1 are included in Table 2. We also provide comprehensive results in Appendix C.2 including for the LLAMA-2 70B Chat and LLAMA-3 70B Instruct models (Table 8) and Mistral Large 2 (Table 9). We find that LM Prompting always achieves the best literal theory of mind performance. Adding the QA format seems to make literal theory of mind a lot worse by taking the input distribution further from the LLM task and closer to the kind of processing needed for actual decision making. The gap between functional theory of mind and literal theory of mind performance is predictably smaller in this case. CoT Prompting has an inconsistent effect on literal theory of mind and always appears to make the functional theory of mind performance better than vanilla QA Prompting. Reflexion and Plans+Insights also seem to improve on QA prompting in terms of functional theory of mind performance, although we find that they do not quite match the performance of generic CoT prompting.
@@ 4.2 Reasoning Over Long Contexts
Deeper Analysis. We have conducted a number of experiments that we do not have space to recap in detail within the main text. In Appendix C.3 we validate the generality of our findings regarding functional theory of mind performance across open source models including LLAMA-3 70B Instruct, Mixtral 8x7B Instruct v1, and Mistral Large 2 when paired with more dynamic tit-for-tat style partner policies. Additionally, in Appendix C.4 we look at the role of inductive bias in the action representation on performance of LLAMA-3 70B Instruct and Mistral Large 2 models. We find that more inductive bias is helpful with a small number of interactions, but could lead to poor long-term behavior that does not converge with respect to the partner’s policy.
Reasoning Models. As our results suggest that achieving functional theory of mind seems related to reasoning over long interaction histories, it is natural to question how the recent emergence of trained chain of thought reasoning models with verifiable rewards such as DeepSeek-R1 (Shao et al., 2024) has impacted theory of mind in these models. In Table 4 we compare the performance of the DeepSeek-R1-Distill-Qwen-32B with the Tabular RMax baseline when using the canonical action names for each game. We do indeed see the strongest functional theory of mind performance that we have seen from an LLM thus far when playing with single action partners. It is competitive with or even exceeding the performance of RMax in all games with single action partners. That said, performance is not as consistent with tit-for-tat style dynamic partners. Notably, we also see a bizarre and unique trend in which a rational response to its literal theory of mind predictions is consistently much poorer than the actual behavior generated in the cases when the functional theory of mind performance is impressive.
These results serve to strengthen our case that functional theory of mind and literal theory of mind performance do not directly imply each other. They also provoke interesting questions about the prediction capabilities that may be lost during the training of reasoning models.
| Model | Game | Partner | ToM % | ||
| Tabular | RPS | Single Action | 0.083 0.004 | 0.039 0.006 | 97.4 0.4 |
| DeepSeek-R1 Distilled 32B | RPS | Single Action | 0.074 0.009 | 0.544 0.126 | 63.7 8.4 |
| Tabular | RPS | Tit-For-Tat | 0.224 0.007 | 0.105 0.000 | 93.0 0.0 |
| DeepSeek-R1 Distilled 32B | RPS | Tit-For-Tat | 0.906 0.041 | 0.921 0.029 | 38.6 2.0 |
| Tabular | IBS | Single Action | 0.211 0.012 | 0.088 0.020 | 98.7 0.3 |
| DeepSeek-R1 Distilled 32B | IBS | Single Action | 0.126 0.045 | 0.233 0.078 | 97.2 0.8 |
| Tabular | IBS | Tit-For-Tat | 0.468 0.031 | 0.162 0.005 | 98.1 0.1 |
| DeepSeek-R1 Distilled 32B | IBS | Tit-For-Tat | 0.045 0.011 | 0.466 0.065 | 94.5 0.8 |
| Tabular | IPD | Single Action | 0.086 0.009 | 0.071 0.015 | 98.6 0.3 |
| DeepSeek-R1 Distilled 32B | IPD | Single Action | 0.121 0.014 | 0.360 0.111 | 83.1 5.8 |
| Tabular | IPD | Tit-For-Tat | 0.248 0.005 | 0.070 0.000 | 98.0 0.0 |
| DeepSeek-R1 Distilled 32B | IPD | Tit-For-Tat | 4.789 0.061 | 0.945 0.139 | 88.9 1.6 |
@@ 5 Alternative Views
We should embrace game theory, not theory of mind. This is perhaps the most prominent alternative view to our paper that was already highlighted in our Rock, Paper, Scissors example. Many papers in the multi-agent RL literature consider finding a game theoretic equilibrium solution as their ultimate goal (Littman et al., 2001; Wang & Sandholm, 2002; Greenwald et al., 2003; Zinkevich et al., 2005, 2007). In some ways game theoretic solutions are opposite to solutions that are found with theory of mind because they focus on performance against worst case or optimal partners rather than the partners agents they are actually paired with. Likewise, they focus on eliminating the exploitability of solutions by other agents as opposed to maximizing expected return in the context of the actual policies of these agents. Proponents of the game theoretic approach will readily acknowledge that theory of mind is more descriptive of the way that humans tend to act (Colman, 2003; Larson, 2004; Owen, 2013). The argument is rather that game theory is prescriptive of the way that AI agents should ideally act. We believe that there is indeed some validity to this point and acknowledge that in a closed system in which all involved agents are AI, this is the optimal case. However, this does not hold as the community moves towards increasingly real-world use cases in which AI must interact with humans or AI agents that behave sub-optimally. These use cases are the norm for LLMs and are the setting in which researchers are interested in measuring LLM theory of mind capabilities. Here there is a gap between the expected reward achievable with the theory of mind solution that optimizes for the expected case and the game theoretic solution that optimizes for the worst case. Moreover, this gap should only get larger as the agent interacts with more humans. We definitely acknowledge that what society will end up preferring depends on the alignment of incentive structures. Humans will probably prefer theory of mind solutions when their relationship with the agent is mostly collaborative, but will not want to be exploited when their relationship with the agent is mostly adversarial. We strongly support instituting guardrails in these cases and believe that this makes it even more important to measure functional theory of mind performance to understand the extent of exploitation by AI models. We provide further detail about the connection between functional theory of mind and relevant equilibria concepts in Appendix B.
@@
Literal theory of mind is paramount for insights related applications. For some applications, we only desire that AI agents provide us with insights, which we will use to make decisions. For these applications, we do not wish to provide the AI with the agency to make these decisions for us. There are indeed very important use cases of this type for which literal theory of mind would be a sufficient capability without functional theory of mind. Generally, these applications are most helpful when the AI has access to more information than the human pertaining to the situation. However, often times the amount of agency we decide to give to an AI depends on its expected reliability in making actual decisions. In this paper we argue that functional theory of mind serves as a necessary test for deciding if AI is ready to receive this agency and that literal theory of mind on its own is not reliable for assessing this performance. It is also important to note that in most applications where we do not provide AI with the agency to act, it would still be useful to humans to have the AI provide an action recommendation.
@@
The word “broken” is too strong. It can be argued that because there are many ways to measure literal theory of mind, adding functional theory of mind is a just single additional aspect within a comprehensive theory of mind evaluation. However, we think “broken” is an appropriate term here because functional theory of mind is what we really care about for applications where we give the AI agency to act, and we have found that deploying an LLM in this setting simply based on its literal theory of mind performance can be misleading to a potentially dangerous extent.
@@
Functional theory of mind is easy with fine-tuning. RL style fine-tuning would easily solve the functional theory of mind problem displayed in our matrix game based experiments as highlighted with simple tabular models. However, our experiments still demonstrate a case where LLM literal theory of mind capabilities do not imply functional theory of mind capabilities and thus highlight the importance of directly measuring functional theory of mind. Most prominent LLM based services today also do not actually tune the model weights separately for each user and rely on in-context learning. The likely reason for this is partly about hardware related restrictions and partly about the difficulty of performing reliable continual learning. See Appendix B for a detailed discussion of connections to continual learning.
The “predict-then-optimize” perspective. Our paper is also related to the end-to-end predict-and-optimize and two-stage predict-then-optimize literature (Elmachtoub & Grigas, 2022). End-to-end approaches that jointly optimize prediction and decision-making often outperform two-stage approaches where prediction is performed independently from optimization because not all prediction errors are equally consequential for the final decision quality. This perspective nicely highlights how analyzing literal theory of mind can lead to misleading conclusions about true functional theory of mind performance, even with ideal self-consistent models.
@@ 6 Conclusion
In this paper, we have argued that existing theory of mind benchmarks are broken because they measure literal theory of mind (Definition 2.1), and not functional theory of mind (Definition 2.2), which is what we typically actually care about in multi-agent social contexts. In Section 4 and Appendix C we provided comprehensive evidence that functional theory of mind performance can be quite poor even when literal theory of mind performance is good and vice versa. Our analysis considers LLMs playing simple matrix games with very simple partners to underscore this point. This position paper serves as a call to action to develop suitably complex LLM theory of mind benchmarks that directly measure functional theory of mind when acting with partners from different personas and social contexts.
Acknowledgements
MR acknowledges support from the Canada Excellence Research Chairs (CERC) Program. We are grateful for the valuable feedback we received from the ICML 2025 reviewers, which led to substantial improvements to our manuscript. Finally, we would like to thank Tim Klinger and Francesca Rossi for their valuable feedback regarding the use of the term “theory of mind” in other fields that helped shape the way we position our view in Section 1.
Appendix A Literal Theory of Mind and Functional Theory of Mind in Popular Theory of Mind Games
In this section, we provide an explanation of how literal theory of mind and functional theory of mind materialize in popular theory of mind focused games. We refer the reader to the cited papers for full descriptions of how each game is played.
Codenames (Bills et al., 2025). In Codenames, literal theory of mind could naturally be measured in terms of guesses by the guesser about the board-card assignments. These assignment can be retrieved from some transformation of the actions that the clue giver would take in different game states following Definition 2.1. Functional theory of mind would be the guesser’s regret with respect to the actual team performance on the game following Definition 2.2, which also includes an aspect of reasoning on top of a literal theory of mind model. For example, given the context of the game, guessers may decide to save certain guesses for later to avoid the potential negative consequences of guessing wrong and potentially gifting the other team an opportunity. We argue in this paper, that it is quite possible that a Codenames agent could perform well in terms of literal theory of mind on this kind of theory of mind task while simultaneously struggling to turn that into a rational strategy, displaying poor functional theory of mind as a result.
Hanabi (Bard et al., 2020). In Hanabi, literal theory of mind could be measured in terms of each agent’s belief about what their own cards are. This belief can be retrieved from some transformation of the actions that other agents took in different game states following Definition 2.1. Functional theory of mind would be the regret with respect to the team score actually achieved at the end of the game. Thus achieving functional theory of mind requires reasoning on top of the theory of mind model regarding how the agent’s cards effects the best actions for the agent to take at various points in the game. It is once again quite possible that a Hanabi agent could perform well in terms of literal theory of mind on this kind of theory of mind task while simultaneously struggling to turn that into a rational strategy in response.
Taboo (Ashktorab et al., 2020). In Taboo, literal theory of mind could be measured in terms of the clue giver’s understanding of the strength of word associations for the particular guesser they are playing with. These associations can be retrieved from some transformation of the actions that the guesser has taken in different previous situations following Definition 2.1 (inside of or outside of the context of the game). Functional theory of mind would be the regret with respect to the team score actually achieved at the end of the game. This again requires additional reasoning on top of the literal theory of mind model, particularly because of the banned list of words that the clue giver cannot say. Thus the clue giver may be inefficient in their search through the space of association strengths even if they have a strong literal theory of mind model.
Wavelength (Morrison et al., 2025). In Wavelength, literal theory of mind could be measured similarly to Taboo in terms of understanding the strength of word associations for the particular partner they are playing with. These associations can be retrieved from some transformation of the actions that the guesser has taken in different previous situations following Definition 2.1 (inside of or outside of the context of the game). Functional theory of mind would again be the regret with respect to the team score actually achieved at the end of the game. The additional reasoning required in this case for achieving functional theory of mind is an understanding of how the other agent thinks about the spectrum dial calibration. Thus the agent can still guess wrong regarding the spectrum even if have a perfect understanding of the other agent’s associations regarding the clue phrase.
@@ Appendix A Literal Theory of Mind and Functional Theory of Mind in Popular Theory of Mind Games
Appendix B Connections to Equilibria Concepts, Continual Learning, and Potential Applications
@@
In-Context Learning vs. Continual Learning. In a multi-agent environment, the environment is nonstationary from the perspective of each agent if the policies of the other agents it interacts with are changing or learning over time (Littman, 1994). As a result, the setting of rapid adaptation to new agents in our paper can be considered a special case of continual RL (see Proposition 3 of Khetarpal et al. (2020)). Proposition 2 of Khetarpal et al. (2020) established that all continual RL problems can be modeled as partial observable problems with agents conditioned on the full interaction history. So, in principle, in-context learning over sufficiently long context representations with sufficiently expressive neural networks can be considered a general solution to continual learning problems. However, it remains to be seen if it is possible to see in-context behaviors resembling continual learning strategies like reservoir sampling replay buffers (Riemer et al., 2018a) or scalable memory efficient approximate buffers (Riemer et al., 2017a, 2019; Bashivan et al., 2019). It would be interesting to also consider policy changes from step to step and their degree of correspondence with older models similar to work leveraging knowledge distillation for continual learning (Li & Hoiem, 2017; Riemer et al., 2017b; Kirkpatrick et al., 2017). Intuitively, chain of thought reasoning may be beneficial because of its similarity with continual learning approaches that can select which layers to process at inference time (Rosenbaum et al., 2018; Cases et al., 2019; Chang et al., 2019; Rosenbaum et al., 2019a), see (Rosenbaum et al., 2019b) for a survey of approaches. More generally, this adaptive computation problem can be formalized within the coagent networks framework (Thomas, 2011; Kostas et al., 2020; Zini et al., 2020). That said, the compositional generalization needed for utilizing these models in practice makes achieving real-world success very challenging (Klinger et al., 2020), potentially making LLM style training a more viable strategy for learning this composition. That said, a word of caution comes from recent theoretical insights about the difficulty of efficiently evaluating models with large context lengths (Riemer et al., 2024a) acting in complex environments (Riemer et al., 2022). This could be a leading reason for the poor performance we are currently seeing with LLMs. Ultimately, there are many potentially interesting continual learning settings that should be considered for a comprehensive evaluation (Normandin et al., 2021), which we leave to future work.
@@
Alternatives to Continual Training. While our work highlights the limitations of using in-context learning for continual learning with current LLMs, full fine-tuning or continual training can be quite expensive in the context of LLMs – especially when it must be done for each end user. However, more computationally efficient alternatives like parameter efficient tuning (Hu et al., 2021; Dettmers et al., 2024; Thakkar et al., 2024a) or model merging (Ilharco et al., 2022; Yadav et al., 2024; Akiba et al., 2024; Thakkar et al., 2024b) may constitute a more economical middle ground. We leave exploration of the comparative efficacy of these strategies to future work. It is important to note that neural scaling laws only demonstrate generalization improvements with bigger model sizes when there is access to as much data as needed (Kaplan et al., 2020). In fact, capacity limits have been found to enhance generalization for RL (Malloy et al., 2020b, 2023) and multi-agent RL (Malloy et al., 2020a, 2021b, 2021a) in the limited data regime. Smaller models also may be necessary to maintain performance in realtime environments that are sufficiently stochastic (Riemer et al., 2024c, b). As such, it may make more sense to consider smaller LLMs if we plan on doing user specific finetuning or auxiliary objectives we would like to maintain like moral values (Padhi et al., 2024; Dognin et al., 2024).
@@
Appendix C Experiment Details
For our experiments we ran each LLM model using the Watsonx API playing a sequence of 100 step episodes with selected policies for the other agents over the course of 24 hours or a maximum of 100 episodes. Every model was evaluated for at least 30 episodes and confidence intervals are based on the actual sample size considered for each LLM model. The payoff tables used for each game are provided in Tables 5, 6, and 7. We also considered the following representations for each action as indicated in the main text: